OpenStudy (jeffthekiller):

what is 782/27 simplified?

OpenStudy (tkhunny):

Have you considered prime factorization? \(\dfrac{782}{27} = \dfrac{391*2}{3*9}\) Can we do more? What does "simplified" mean? Are we seeking a Mixed Fraction?

OpenStudy (jesselacompte):

27

OpenStudy (jesselacompte):

8

OpenStudy (jeffthekiller):

i mean i need a mixed number, ;(

OpenStudy (jeffthekiller):

@jesselacompte there is an answer of 28 and 26/27....

OpenStudy (tkhunny):

This may require division. 782/27 = 28.962962962962962962962962962963 That does narrow it down quite a bit.